Consider a section given below. Given that c = 11 mm. Determine I(x1) and I(x2) and I(y) in mm^4.
|-6 mm-f-6 mm-
Answer #1

moment of inertia of rectangular body about centroidal x axis

  (INA), =

moment of inertia of rectangular body about centroidal y axis

  (INA)y = 12

moment of inertia of triangular body about centroidal x axis

(INA): = 36

moment of inertia of triangular body about centroidal y axis

(INA); = 36

parallel axis theorem

I = INA + AY

Controidal raxis s R--- Xg C =//mm Controidal X-axis Na -coala . Kemm ² mm for section O ADE Area = A= OAXAD = 6X11= 66 mm2 ddistance af controld af ABD from y-axis =&g = OA+ AB - 67 - 8mm from x,-axis .: Centroid of cross-section 5 = Axt Az xg - Alt(INA), = 961.2856 mmt moment of inertia about cont voidal Y-axis (Java) = ADX 903 47 (2,-, JP + ADX PAR (79-7 )? Ins) y = 1moment af Inertial about Y-axis Iy = (INA), + A la Py = 814.0011 +998 4.674= 2973.0822 mmt (Ans